# Stellar reddening map from DESI imaging and spectroscopy

### v1

@@ Abstract
We present new Galactic reddening maps of the high Galactic latitude sky using DESI imaging and spectroscopy. We directly measure the reddening of 2.6 million stars by comparing the observed stellar colors in $g-r$ and $r-z$ from DESI imaging with the synthetic colors derived from DESI spectra from the first two years of the survey. The reddening in the two colors is on average consistent with the [Fitzpatrick (1999)] extinction curve with $R_{\mathrm{V}}=3.1$. We find that our reddening maps differ significantly from the commonly used [Schlegel et al. (1998)] (SFD) reddening map (by up to 80 mmag in $E(B-V)$), and we attribute most of this difference to systematic errors in the SFD map. To validate the reddening map, we select a galaxy sample with extinction correction based on our reddening map, and this yields significantly better uniformity than the SFD extinction correction. Finally, we discuss the potential systematic errors in the DESI reddening measurements, including the photometric calibration errors that are the limiting factor on our accuracy. The $E(g-r)$ and $E(g-r)$ maps presented in this work, and for convenience their corresponding $E(B-V)$ maps with SFD calibration, are publicly available.
@@ 1 Introduction
Galactic dust absorbs and scatters light mostly in the optical and re-emits in the infrared as thermal radiation. The dust attenuation in the optical wavelengths can be a significant nuisance for galaxy surveys as it makes sources behind the dust appear dimmer and redder. The intrinsic brightness and color of an extragalactic source are usually inferred by applying a correction to the observed brightness and color. The amplitude of this correction is often calculated using a reddening map that provides the reddening for a specific wavelength or color and an extinction curve (or “extinction law”) that prescribes the amount of extinction as a function of wavelength.
@@ 2 Data
We use the DESI spectroscopic data from the first two years of observations, from May 14, 2021, to June 21, 2023, to obtain the synthetic stellar spectra and the resulting synthetic colors. We use the Y1 reduction (internally named the “iron” data release) for the first-year data (taken from May 14, 2021, to June 14, 2022) and the internal “daily” reduction of the second-year data (taken from the end of Y1 to June 21, 2023). Specifically, for the first-year data we use the HEALPix coadds (which co-add across overlapping tiles; see §3.3.2 of DESI Collaboration et al. 2024), and for the second-year data we use the tile-based coadds (in which repeat observations in overlapping tiles are not combined).
@@
    Standard stars in all three programs (STD_FAINT and STD_BRIGHT in dark and bright programs; GAIA_STD_FAINT and GAIA_STD_BRIGHT in the backup program);
@@
    Milky Way Survey (MWS) magnitude limited blue sample, red sample, and bulk sample in the bright program (MWS_MAIN_BLUE, MWS_MAIN_RED, and MWS_BROAD);
@@
    Brighter and fainter backup targets in the backup program (BACKUP_BRIGHT and BACKUP_FAINT).
@@
Table 1 lists the number of stars from each target type in the final sample. Figure 1 shows the density map of the final stellar sample. The total sky coverage is 13,835 square degrees, based on the number of HEALPix pixels with NSIDE=256 that contain at least one star in the $g-r$ reddening map.
@@
Table 1: Statistics of the different types of stars used for the reddening measurement, i.e., after the selection cuts (see text for details). The number of stars ($N_{g-r}$ and $N_{r-z}$) is the total number of stars used in the $g-r$ and $r-z$ reddening measurements (after passing the quality cuts). There are 2205615 stars that pass both the $g-r$ and $r-z$ quality cuts. Note that there is overlap between some of the target classes (e.g., STD_BRIGHT is entirely a subset of STD_FAINT). The $\sigma(E(g-r))$ and $\sigma(E(r-z))$ columns list the per-star error in mmag (millimagnitude) in the reddening measurements; the errors are obtained by comparing the per-star measurements with the map pixel values (which average over the reddening measurement of at least 32 stars); the first number is the root-mean-square (RMS) error, and the second number in parentheses is normalized median absolute deviation (NMAD), here defined as $\sigma_{\mathrm{NMAD}}(\Delta x)=1.4826\times\mathrm{Median}(|\Delta x|)$. Some target classes have much larger RMS errors than NMAD errors because they have a small fraction of stars with large errors (to which $\sigma_{\mathrm{NMAD}}$ is not sensitive).
Figure 1: The density of stars used in $g-r$ reddening measurements (see Section 3). The gaps and missing pixels are due to the lack of coverage from the first two years of DESI observations. The curve separating the two regions is the Galactic plane. The HEALPix resolution is NSIDE=256 (13.7 pixel size).
@@ 2.1 Modeling the stellar spectra: rvspecfit
We model the observed spectra using the template fitting code rvspecfit (Koposov, 2019; Koposov et al., 2011) to obtain the synthetic colors. rvspecfit was designed to fit non-flux calibrated spectra. Thus it models the observed spectra $S(\lambda)$ by $T(\lambda\sqrt{\frac{1+v/c}{1-v/c}},p)*P(\lambda)$ where $T(\lambda,p)$ is the interpolated stellar spectrum for a given set of stellar parameters $p$, $v$ is the radial velocity, and $P(\lambda)$ is either a polynomial or linear combination of any set of basis functions (such as radial basis functions (RBFs)). For DESI, $P(\lambda)$ currently consists of a 2nd order polynomial (1, $\lambda$, $\lambda^{2}$) and 7 Gaussian RBFs centered at equally spaced points along wavelength. The standard deviation of each Gaussian RBF is 1/7 of the wavelength range. (For simplicity, we refer to $P(\lambda)$ as “the polynomial term”.) The grid of stellar spectra used by rvspecfit is the PHOENIX v2.0 grid by Husser et al. (2013) convolved to the average DESI resolution of FWHM=1.55 Å$\mathrm{FWHM}=1.55\textup{~{}\AA}$ for blue and red cameras and FWHM=1.8 Å$\mathrm{FWHM}=1.8\textup{~{}\AA}$ for the near-infrared camera.
@@
For each star rvspecfit produces a best-fit synthetic spectrum with zero extinction as well as the following stellar atmosphere parameters that determine the synthetic spectrum: effective temperature $T_{\mathrm{eff}}$, surface gravity log $g$, metallicity [Fe/H], and alpha element abundance [$\alpha$/Fe]. Figure 2 shows an example DESI spectrum and the rvspecfit best-fit model. We use the best-fit spectral models without the polynomial term to compute zero-extinction synthetic colors and compare them with the observed colors from imaging to determine the Galactic reddening. Given this objective, it is important to note that because of the multiplicative polynomial term $P(\lambda)$, the stellar parameters mostly only depend on the narrow absorption features, and the actual amount of extinction has little impact on the best-fit model (at least within the maximum amount of extinction that we are concerned with). We will discuss this more in Section 5.3. This polynomial term removes the dependence of not only extinction but also other systematic errors such as flux calibration errors in DESI spectra.
@@ 2.2 Quality cuts
We first remove stars within 1 kpc from us by requiring that the Gaia parallax be less than 1 mas (milliarcsecond). The DESI footprint is almost entirely above the galactic latitude of $|b|=18^{\circ}$; at this galactic latitude, a distance of 1 kpc corresponds to a scale height of 0.31 kpc. The parallax cut ensures that we are always measuring the total dust extinction at any line of sight. The parallax cut removes less than 2% of the stars.
@@
We also try to identify and remove photometric outliers by comparing with Gaia photometry: we use Gaia magnitudes in G, BP and RP bands to predict magnitudes in LS filters, and we remove any stars whose observed magnitude differs from the predicted magnitude by more than 0.1 mag. This removes less than 1% of the stars. Details of the Gaia-to-LS transformation are described in Appendix A.
We also apply spectroscopic quality cuts to remove spectra with low S/N or flagged by the DESI pipeline (Guy et al., 2023), specifically the redshift fitter redrock [^2] https://github.com/desihub/redrock (Bailey et al., 2024). We require that the object is classified as a star, and that it is not a white dwarf (which are not included in the PHOENIX synthetic spectra and thus cannot be accurately modeled by rvspecfit).
Finally, we select stars within a certain range of stellar atmosphere parameters. The limits on the stellar parameters are decided empirically based on the accuracy of the synthetic colors from the rvspecfit model. Stars with parameters outside the range typically have significantly larger errors in their synthetic color than those within the range; see the upper panel of Figure 3. The stellar parameter cuts remove roughly 23% of the stars (most of the removed stars are below the $T_{\mathrm{eff}}$ threshold of 5000 K). The selected temperature range of 5000–6500 K roughly corresponds to G and F spectral types.
@@ 3.1 Method
If we assume a universal extinction curve, the amount of extinction at some wavelength for a given line of sight can be expressed as the wavelength-dependent extinction coefficient $R(\lambda)$ multiplied by the reddening $E$ in a pre-defined color at that line of sight
@@
For this work, we assume the extinction curve of Fitzpatrick (1999) with $R_{\mathrm{V}}=3.1$. The extinction coefficients $R_{x}$ are computed based on an F dwarf star with 7000 K and [Fe/H]=$-1$. Table 3 lists the extinction coefficients for our filters. The table also lists the effective wavelengths, which are defined as the wavelength where the extinction coefficient equals that of the filter, i.e., $R(\lambda_{\mathrm{eff}})=R_{x}$. (This definition is allowed because the extinction curve is monotonic in the optical wavelengths and thus invertible.) The extinction coefficients are slightly different for the Northern and Southern imaging due to differences in the filter curves and (to a lesser extent) in the average airmass. For the DECam filters (Southern imaging), we assume an airmass of 1.3. For the BASS and MzLS filters (Northern imaging), we assume an airmass of 1.1.
Since it is inconvenient to keep track of two slightly different filter sets for the reddening maps, we renormalize the Northern reddening measurements to match the $R_{a}-R_{b}$ value of the South before producing the reddening maps and only report the reddening as would be measured the DECam filters (see 3.4).
@@ 3.2 Calibration of synthetic colors
With the initial offset correction, we can now compute the average reddening for stars in each HEALPix pixel and create a HEALPix map with NSIDE of 128 (pixel size of 27.4). Only pixels containing at least 16 stars are used. We repeat the procedure to calculate the per-star offset, now obtained by comparing the per-star reddening of each of the 400 nearest neighbors in stellar parameter space with the mean reddening of its corresponding pixel, and update the HEALPix map with the updated offset. The per-star offset calculation and map update are iterated for a total of 6 times to obtain the final offset. The procedure is performed separately for $g-r$ and $r-z$.
@@ 3.3 Error estimation of reddening measurements
We first compute the difference between the per-star reddening measurement $E(a-b)_{\mathrm{star}}$ and the average reddening of the corresponding HEALPix pixel (with NSIDE of 128) $E(a-b)_{\mathrm{map}}$ whose measurement error is much smaller than the individual measurement. We treat this difference as a result of the uncertainties in the reddening measurements. We only use stars in pixels with at least 32 stars and with $E(B-V)_{\mathrm{SFD}}<0.1$. We then divide the sample into 2-D bins of SN_B and $\sigma_{a-b,\mathrm{NN}}$, keeping only bins with at least 16 stars, and for each bin compute $\sigma_{a-b}$, which is the RMS in $E(a-b)_{\mathrm{star}}-E(a-b)_{\mathrm{map}}$. We fit $\sigma_{a-b}$ as a function of SN_B and $\sigma_{a-b,\mathrm{NN}}$ with 5th-order polynomials using the statsmodels’ Robust Linear Models (RLM) routine, and compute the per-star reddening error using the polynomials. We impose an error floor of 12 mmag for both $E(g-r)$ and $E(r-z)$. Figure 6 shows the measured per-bin RMS errors, the model, and the residuals for the bright-time MWS stars in the South. For stars belonging to bins with fewer than 16 stars (typically in outlying regions in SN_B and $\sigma_{a-b,\mathrm{NN}}$ where the polynomials may not be well behaved), they are assigned the error of their nearest neighbor in normalized SN_B - $\sigma_{a-b,\mathrm{NN}}$ space.

### v3

@@ Abstract
We present new Galactic dust reddening maps of the high Galactic latitude sky using DESI imaging and spectroscopy. We directly measure the reddening of 2.6 million stars by comparing the observed stellar colors in $g-r$ and $r-z$ from DESI imaging with the synthetic colors derived from DESI spectra from the first two years of the survey. The reddening in the two colors is on average consistent with the [Fitzpatrick (1999)] extinction curve with $R_{\mathrm{V}}=3.1$. We find that our reddening maps differ significantly from the commonly used [Schlegel et al. (1998)] (SFD) reddening map (by up to 80 mmag in $E(B-V)$), and we attribute most of this difference to systematic errors in the SFD map. To validate the reddening map, we select a galaxy sample with extinction correction based on our reddening map, and this yields significantly better uniformity than the SFD extinction correction. Finally, we discuss the potential systematic errors in the DESI reddening measurements, including the photometric calibration errors that are the limiting factor on our accuracy. The $E(g-r)$ and $E(r-z)$ maps presented in this work, and for convenience their corresponding $E(B-V)$ maps with SFD calibration, are publicly available.
@@ 1 Introduction
Galactic dust absorbs and scatters light mostly in UV and optical wavelengths and re-emits in the infrared as thermal radiation. The dust extinction in the optical wavelengths can be a significant nuisance for galaxy surveys as it makes sources behind the dust appear dimmer and redder. The intrinsic brightness and color of an extragalactic source are usually inferred by applying a correction to the observed brightness and color. The amplitude of this correction is often calculated using a reddening map that provides the reddening for a specific wavelength or color and an extinction curve (or “extinction law”) that prescribes the amount of extinction as a function of wavelength.
@@ 2 Data
We use the DESI spectroscopic data from the first two years of observations, from May 14, 2021, to June 21, 2023, to obtain the synthetic stellar spectra and the resulting synthetic colors. We use the Y1 reduction (internally named the “iron” data release) for the first-year data (taken from May 14, 2021, to June 14, 2022) and the internal “daily” reduction of the second-year data (taken from the end of Y1 to June 21, 2023). Specifically, for the first-year data we use the HEALPix  [^2] https://healpix.sourceforge.io/ coadds (which co-add across overlapping tiles; see §3.3.2 of DESI Collaboration et al. 2024), and for the second-year data we use the tile-based coadds (in which repeat observations in overlapping tiles are not combined).
@@
    Standard stars in all three programs (STD_FAINT and STD_BRIGHT in dark and bright programs; GAIA_STD_FAINT and GAIA_STD_BRIGHT in the backup program); see also Guy et al. (2023) and Myers et al. (2023);
@@
    Milky Way Survey (MWS) magnitude limited blue sample, red sample, and bulk sample in the bright program (MWS_MAIN_BLUE, MWS_MAIN_RED, and MWS_BROAD); see also Cooper et al. (2023);
@@
    Brighter and fainter backup targets in the backup program (BACKUP_BRIGHT and BACKUP_FAINT); see also Dey et al. (2025).
@@
Table 1 lists the number of stars from each target type in the final sample. Figure 1 shows the density map of the final stellar sample. The total sky coverage is 13,835 square degrees, based on the number of HEALPix pixels with NSIDE=256 (13.7 pixel size) that contain at least one star in the $g-r$ reddening map.
@@
Table 1: Statistics of the different types of stars used for the reddening measurement, i.e., after the selection cuts (see text for details). The number of stars ($N_{g-r}$ and $N_{r-z}$) is the total number of stars used in the $g-r$ and $r-z$ reddening measurements (after passing the quality cuts). There are 2205615 stars that pass both the $g-r$ and $r-z$ quality cuts. Note that there is overlap between some of the target classes (e.g., STD_BRIGHT is entirely a subset of STD_FAINT). The $\sigma(E(g-r))$ and $\sigma(E(r-z))$ columns list the per-star error in mmag (millimagnitude) in the reddening measurements; the errors are obtained by comparing the per-star measurements with the map pixel values (which average over the reddening measurement of at least 32 stars) of NSIDE=256 (13.7 pixel size) maps; the first number is the root-mean-square (RMS) error, and the second number in parentheses is normalized median absolute deviation (NMAD), here defined as $\sigma_{\mathrm{NMAD}}(\Delta x)=1.4826\times\mathrm{Median}(|\Delta x|)$. Some target classes have much larger RMS errors than NMAD errors because they have a small fraction of stars with large errors (to which $\sigma_{\mathrm{NMAD}}$ is not sensitive).
Figure 1: The density of stars used in $g-r$ reddening measurements (see Section 3). The gaps and missing pixels are due to the lack of coverage from the first two years of DESI observations. The curve separating the two regions is the Galactic plane. The HEALPix resolution is NSIDE=256 (13.7 pixel size). The color scale is linear in this figure and all other HEALPix map figures.
@@ 2.1 Modeling the stellar spectra: rvspecfit
We model the observed spectra using the template fitting code rvspecfit (Koposov, 2019; Koposov et al., 2011) to obtain the synthetic colors. rvspecfit was designed to fit non-flux calibrated spectra. Thus it models the observed spectra $S(\lambda)$ by $T(\lambda\sqrt{\frac{1+v/c}{1-v/c}},p)*P(\lambda)$ where $T(\lambda,p)$ is the interpolated stellar spectrum for a given set of stellar parameters $p$, $v$ is the radial velocity, and $P(\lambda)$ is either a polynomial or linear combination of any set of basis functions (such as radial basis functions (RBFs)). For DESI, $P(\lambda)$ currently consists of a 2nd order polynomial (1, $\lambda$, $\lambda^{2}$) and 7 Gaussian RBFs centered at equally spaced points along wavelength. The standard deviation of each Gaussian RBF is 1/7 of the wavelength range.
Since $P(\lambda)$ is a linear combination of fixed functions, the best values of the linear coefficients (in a $\chi^{2}$ sense) for a given $P(\lambda)$ can be found through a simple matrix operation (see Koposov et al. 2011 for more details). Thus, when fitting the stellar spectra, the optimization is performed over $P(\lambda)$ and radial velocity using a combination of Nelder-Mead and BFGS optimizers (see also Cooper et al. 2023). (For simplicity, we refer to $P(\lambda)$ as “the polynomial term”.)
The grid of stellar spectra used by rvspecfit is the PHOENIX v2.0 grid by Husser et al. (2013) convolved to the average DESI resolution of FWHM=1.55 Å$\mathrm{FWHM}=1.55\textup{~{}\AA}$ for blue and red cameras and FWHM=1.8 Å$\mathrm{FWHM}=1.8\textup{~{}\AA}$ for the near-infrared camera.
@@
For each star rvspecfit produces a best-fit synthetic spectrum with zero extinction as well as the following stellar atmosphere parameters that determine the synthetic spectrum: effective temperature $T_{\mathrm{eff}}$, surface gravity log $g$, metallicity [Fe/H], and alpha element abundance [$\alpha$/Fe]. Figure 2 shows an example DESI spectrum and the rvspecfit best-fit model. We use the best-fit spectral models without the polynomial term to compute zero-extinction synthetic colors and compare them with the observed colors from imaging to determine the Galactic reddening. Given this objective, it is important to note that because of the multiplicative polynomial term $P(\lambda)$, the stellar parameters mostly only depend on the narrow absorption features, and the actual amount of extinction has little impact on the best-fit model (at least for the typical amount of extinction, $E(B-V)\lesssim 0.2$, encountered by most cosmological probes). We will discuss this more in Section 5.3. This polynomial term removes the dependence of not only extinction but also other systematic errors such as flux calibration errors in DESI spectra.
@@ 2.2 Quality cuts
We first remove stars within 1 kpc from us by requiring that the Gaia parallax be less than 1 mas (milliarcsecond). The DESI footprint is almost entirely above the galactic latitude of $|b|=18^{\circ}$. This corresponds to a distance of 310 pc from the Galactic plane at the lowest latitude in the survey, which is significantly larger than the typical dust scale height of $\sim$100 pc, ensuring that the stars used are behind most of the dust. While the 1 kpc limit should be adequate for most of the high Galactic latitude sky, we will investigate more optimized parallax cuts in future studies.
@@
We also try to identify and remove photometric outliers by comparing with Gaia photometry: we use Gaia magnitudes in G, BP and RP bands to predict magnitudes in LS filters, and we remove any stars whose observed magnitude differs from the predicted magnitude by more than 0.1 mag. This removes less than 1% of the stars. Details of the Gaia-to-LS transformation are described in Appendix A. Roughly 40% of the stars removed by this cut have Gaia Renormalized Unit Weight Error (RUWE) [^3] Document GAIA-C3-TN-LU-LL-124-01 from https://www.cosmos.esa.int/web/gaia/public-dpac-documents larger than 1.4. For comparison, less than 1% of the remaining stars have RUWE$>$1.4. Gaia’s recommended criterion of good astrometric solutions of single stars is RUWE$<$1.4. Therefore, many (if not most) of the stars removed by the Gaia photometric outlier cut may be double stars.
We also apply spectroscopic quality cuts to remove spectra with low S/N or flagged by the DESI pipeline (Guy et al., 2023), specifically the redshift fitter redrock [^4] https://github.com/desihub/redrock (Bailey et al., 2025). We require that the object is classified as a star, and that it is not a white dwarf (which are not included in the PHOENIX synthetic spectra and thus cannot be accurately modeled by rvspecfit).
Finally, we select stars within a certain range of stellar atmosphere parameters. The limits on the stellar parameters are decided empirically based on the accuracy of the synthetic colors from the rvspecfit model. Stars with parameters outside the range typically have significantly larger errors in their synthetic color than those within the range; see the upper panel of Figure 3. The stellar parameter cuts remove roughly 23% of the stars (most of the removed stars are below the $T_{\mathrm{eff}}$ threshold of 5000 K). The selected temperature range of 5000–6500 K roughly corresponds to G and F spectral types. While the stars removed by the stellar parameter cuts have larger errors and/or systematic offsets, some of them may still be useful for reddening measurements, and we will explore the broader stellar parameter space in future studies.
@@ 3.1 Method
If we assume an extinction curve, the amount of extinction at some wavelength for a given line of sight can be expressed as the wavelength-dependent extinction coefficient $R(\lambda)$ multiplied by the reddening $E$ in a pre-defined color at that line of sight
@@
To convert our reddening measurements to $E(B-V)$, we assume a universal extinction curve of Fitzpatrick (1999) with $R_{\mathrm{V}}=3.1$. The extinction coefficients $R_{x}$ are computed based on an F dwarf star with 7000 K and [Fe/H]=$-1$. Table 3 lists the extinction coefficients for our filters. The table also lists the effective wavelengths, which are defined as the wavelength where the extinction coefficient equals that of the filter, i.e., $R(\lambda_{\mathrm{eff}})=R_{x}$. (This definition is allowed because the extinction curve is monotonic in the optical wavelengths and thus invertible.) The extinction coefficients are slightly different for the Northern and Southern imaging due to differences in the filter curves and (to a lesser extent) in the average airmass. For the DECam filters (Southern imaging), we assume an airmass of 1.3. For the BASS and MzLS filters (Northern imaging), we assume an airmass of 1.1.
Since it is inconvenient to keep track of two slightly different filter sets for the reddening maps, we renormalize the Northern reddening measurements to match the $R_{a}-R_{b}$ value of the South before producing the reddening maps and only report the reddening as would be measured the DECam filters (see Section 3.4).
@@
Note that while we assume a universal extinction curve to infer $E(B-V)$, our $E(g-r)$ and $E(r-z)$ measurements are direct measurements and they do not rely on assumptions of the extinction curve (except for a slight dependence in the renormalization the Northern reddening measurements).
@@ 3.2 Calibration of synthetic colors
With the initial offset correction, we can now compute the average reddening for stars in each HEALPix pixel and create a HEALPix map with NSIDE of 128 (pixel size of 27.4). Only pixels containing at least 16 stars are used. We repeat the procedure to calculate the per-star offset, now obtained by comparing the per-star reddening of each of the 400 nearest neighbors in stellar parameter space with the mean reddening of its corresponding pixel, and update the HEALPix map with the updated offset. The per-star offset calculation and map update are iterated for a total of 6 times (which are sufficient for the maps to converge) to obtain the final offset. The procedure is performed separately for $g-r$ and $r-z$.
@@ 3.3 Error estimation of reddening measurements
We first compute the difference between the per-star reddening measurement $E(a-b)_{\mathrm{star}}$ and the average reddening of the corresponding HEALPix pixel (with NSIDE of 128) $E(a-b)_{\mathrm{map}}$ whose measurement error is much smaller than the individual measurement. We treat this difference as a result of the uncertainties in the reddening measurements. We only use stars in pixels with at least 32 stars and with $E(B-V)_{\mathrm{SFD}}<0.1$. We then divide the sample into 2-D bins of SN_B and $\sigma_{a-b,\mathrm{NN}}$, keeping only bins with at least 16 stars, and for each bin compute $\sigma_{a-b}$, which is the RMS in $E(a-b)_{\mathrm{star}}-E(a-b)_{\mathrm{map}}$. We fit $\sigma_{a-b}$ as a function of SN_B and $\sigma_{a-b,\mathrm{NN}}$ with 5th-order polynomials using the statsmodels’ Robust Linear Models (RLM) routine, and compute the per-star reddening error using the polynomials. (We choose 5th-order polynomials as they are the lowest order polynomials that can adequately match the observed trends in diagnostic plots such as Figure 6.) We impose an error floor of 12 mmag for both $E(g-r)$ and $E(r-z)$. Figure 6 shows the measured per-bin RMS errors, the model, and the residuals for the bright-time MWS stars in the South. For stars belonging to bins with fewer than 16 stars (typically in outlying regions in SN_B and $\sigma_{a-b,\mathrm{NN}}$ where the polynomials may not be well behaved), they are assigned the error of their nearest neighbor in normalized SN_B - $\sigma_{a-b,\mathrm{NN}}$ space.